OpenStudy (anonymous):

10. Solve the equation. 3m – 4 = 26 (1 point)

OpenStudy (anonymous):

Simplifying 3m + -4 =26

OpenStudy (anonymous):

super easy! 3m - 4 = 26 3m -4 +4 = 26 + 4 3m = 30 3m/3 = 30/3 m= 10 :D good luckkk

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

the sum would be 2.9 but to make sure try and do the work